### v1

@@ 3.3 Finite Mixture Model Application
In a similar context of this paper, mixture models are applied to count data in White and Murphy (2016) to compare the results with those achieved with the mixed membership model. For a more in-depth explanation of the mixture modelling framework please refer to Everitt and Hand (1981) and Bouveyron et al (2019).
We fitted the multivariate Poisson mixture model to the player data and the profiles means are presented in Table 5 and Figure 4. In this application, similar to the MM model, the mixture model encounters challenges in effectively capturing the archetypal players. It’s important to clarify, as highlighted in the Introduction (see Section 1), that the finite mixture models function differently compared to mixed and partial membership models. Specifically, in finite mixture models, the membership of each unit to the clusters is binary, rather than continuous between 0 and 1, as is the case in mixed and partial membership models. This distinction necessitates a different interpretation of the results. In finite mixture models, the posterior probability of cluster membership should be viewed as reflecting uncertainty rather than partial membership. Nonetheless, for a comprehensive understanding, we include the results derived from the finite mixture model.
The players with highest membership per profile are: the attacking midfielder Filip Duricić with 0.570 in profile 1, the full-back Giovanni Di Lorenzo with 0.574 membership in profile 2, the midfielder Filip Kostić with 0.620 membership in profile 3, the defensive player Danilo Luiz da Silva with 0.605 membership in profile 4, the midfielder Stefano Sensi with 0.609 membership in profile 5 and the full-back Rogério Oliveira da Silva with 0.609 membership in profile 6.
Table 6: MM model expected profiles means
Gls Ast PrgC PrgP Sh SoT KP PiFT
1 7.85 2.94 29.21 120.45 41.50 18.24 40.96 41.33
2 2.98 3.44 46.24 78.69 24.83 9.03 25.05 23.75
3 0.75 1.29 7.79 55.00 9.23 3.49 15.50 103.26
4 0.08 0.14 72.16 0.19 64.54 0.09 0.35 11.53
5 0.51 0.27 0.14 37.22 0.22 1.72 3.22 62.41
6 0.77 16.43 18.52 11.94 3.58 8.15 1.37 5.72
PPA CrsPA SCA PassLive PassDead TO ShToSh Def
1 13.31 2.10 1.75 32.32 0.31 4.69 6.00 0.55
2 28.90 12.49 70.33 50.62 5.83 3.65 3.47 0.96
3 14.42 4.53 39.42 29.15 1.39 1.10 1.37 0.69
4 0.50 0.09 3.83 2.47 0.69 0.09 0.09 0.14
5 2.55 0.37 10.63 8.56 21.38 0.11 0.47 0.41
6 1.52 21.68 16.23 0.68 0.23 1.38 0.91 2.06
GCA Tkl Blocks Int Clr Err
1 6.41 10.27 9.51 3.62 17.89 0.21
2 7.78 24.28 15.94 11.20 5.41 0.20
3 2.95 34.89 22.23 18.89 33.38 0.42
4 0.21 2.95 0.16 0.48 54.07 1.00
5 1.05 0.90 18.81 19.66 89.60 1.00
6 53.16 32.30 30.60 13.01 0.59 0.05
Figure 5: Finite mixture model. Expected profiles means, conditional on profile membership, with 6 profiles.
It is clear that in the mixture model, the interpretation of the profile means is less immediate than the PM model or even the MM model.
@@ 3.4 Model Comparison
When comparing clustering models, one of the key parameters to consider is the interpretability of the results. This aspect is extensively discussed in Fraley and Raftery (1998) and Forgy (1965). The ability to interpret the clusters and derive meaningful insights from them is crucial in various domains, including sports analysis.
When comparing the models, it becomes evident that the interpretation of the clusters is significantly easier and more accurate in the PM model compared to the MM model. This disparity can be attributed to the inherent differences in the underlying assumptions of the two models.
The PM model, unlike the MM model, does not assume factorization over attributes. This means that each data attribute of a given data point is not assumed to be drawn independently from a mixture distribution based on the membership vector. In contrast, MM models are designed to handle situations where the objects being modeled consist of exchangeable sub-objects.
The lack of such assumptions in the PM model enhances its interpretability in our application. It allows for a more straightforward and intuitive understanding of the clusters, as the model does not impose strict dependencies between the attributes. As a result, the PM model excels in capturing the nuances of different playing positions without forcing the interpretation to conform to specific attribute relationships. The ease and quality of cluster interpretation are superior in the PM model compared to the MM model and even more when compared with the mixture model. The latter, with its crisp clustering approach, falls short of meeting the specific objectives and requirements of our proposed application. Additionally, the PM model demonstrates superior capability in identifying archetypal players, a key aspect in our analysis.
Overall, the PM model’s flexibility, underpinned by its independence from attribute-related assumptions, enables it to more accurately represent the diverse playing positions in football. This is achieved without the burden of imposing restrictive assumptions on the dataset, making it a highly suitable choice for this application.
@@ 4 Conclusions And Future Developments
Partial membership models offer analysts significantly greater flexibility compared to traditional model-based clustering or standard distance-based clustering methods. In our study, we tailored the model specifically for count data and applied it to the analysis of Serie A football players during the 2022/2023 season. Our goal was to estimate the various positions players tend to occupy on the field, in addition to their primary positions, by analyzing their playing styles. We based our analysis on a set of 22 count variables recorded during games, carefully selected to cover the key skills pertinent to each player’s role. This application also allowed us to test the model’s reliability by comparing its results with each player’s actual playing position. Furthermore, when compared with mixed membership and mixture models for count data, the partial membership model yielded more realistic and interpretable results. It excelled particularly in capturing archetypal players, whose distinct characteristics define specific clusters. These archetypes not only aid in clarifying each cluster’s definition but also contribute to the model’s overall stability and interpretability. By serving as benchmarks, they assist in better identifying and understanding the roles of players whose classifications might otherwise be ambiguous. While the partial membership model does require considerable computational resources, we believe its potential applications extend beyond sports analysis, offering valuable insights in fields like social sciences, genetics, natural sciences, and textual analysis. It can address some limitations inherent in mixed membership models. In the realm of sports analysis, our findings could significantly benefit coaches, talent scouts, team managers, and analysts. Utilizing these insights, they can make more informed decisions regarding team strategy, talent acquisition, and statistical research, thereby enhancing both performance and understanding in football.
For future developments, two primary areas present intriguing opportunities for enhancing our model. Firstly, addressing the issue of over-dispersion, which is a common challenge in count data, is of considerable interest. Developing methods to accurately assess and incorporate over-dispersion into the model would enable more precise and reliable predictions, particularly in datasets where variance significantly exceeds the mean. Secondly, adding a temporal dimension to the model opens up another avenue for exploration. By incorporating a temporal aspect, the model could provide a dynamic view of the clusters and allow for more nuanced analyses.
@@ Appendix A Application: Washington DC Bike Data
We calculated the number of times bikes are collected from each of the 660 stations and we modelled these counts using a partial membership model, with the intent to explore the interactions between the bikes stations usage, to improve the allocation of the bikes. Partial membership model suits this type of application because the bikes move between the stations along the day, so the stations usage could vary and their membership could be partial. It should be noted that we do not addressed the temporal dependency as the temporal nature of the data would require. Nevertheless, the approaches appear to identify interesting behaviour in the data, and serve to illustrate the usefulness of the method. We run the model over a range of $K=1,\dots,6$. The model with the lowest WAIC is the one with 5 profiles (or components). For a better visualization, in Figure 7 are represented the natural log of the profiles means, while Figure 8 shows the marginal simplices representing stations’ profile membership.

### v2

@@ 3.3 Finite Mixture Model Results
Mixture models for count data have been applied in contexts similar to ours (e.g., White and Murphy, 2016), often to compare with mixed membership approaches. For a foundational explanation of mixture modeling, we refer to Everitt and Hand (1981) and Bouveyron et al (2019).
We fit a multivariate Poisson mixture model to the player data. The component-specific means are shown in Table 6 and illustrated in Figure 6. Because finite mixture models assign each player to exactly one component, they do not explicitly accommodate players who truly balance multiple roles. Instead, they treat multi-role players as uncertain assignments to a single “best-fitting” component.
Table 6: Finite mixture model: expected component means (Poisson rate parameters).
Gls Ast PrgC Sh KP CrsPA SCA TO
1 7.00 5.44 71.51 55.83 48.87 9.98 97.65 6.12
2 0.98 1.37 29.07 16.93 15.53 3.99 40.32 0.69
3 0.05 0.23 0.05 0.05 0.71 0.05 5.05 0.05
4 1.15 0.55 11.09 11.76 4.69 0.75 15.62 0.23
5 2.43 2.60 61.94 27.44 31.13 9.19 69.52 2.93
6 2.48 35.57 52.53 24.10 1.84 56.84 4.17 5.74
ShToSh Def GCA Tkl Blocks Int Clr
1 6.12 0.90 11.21 30.24 21.40 12.90 13.98
2 1.93 1.03 3.34 55.80 35.20 32.76 60.96
3 0.05 0.09 0.52 1.01 0.15 0.67 12.51
4 1.29 0.53 1.36 37.63 31.98 29.42 93.07
5 3.23 1.30 6.58 50.16 29.04 23.19 35.71
6 0.83 6.43 15.43 13.97 5.45 17.09 0.13
Figure 6: Finite mixture model with 6 components. Poisson means for each variable are plotted for each component.
Based on the most prominent variables, we interpret the six components as follows:
  • Component 1: Demonstrates notably high Goals (7.00), Shots (55.83), Key Passes (48.87), and Shot-Creating Actions (97.65). Though there is also a moderate defensive contribution (e.g., Tkl $=30.24$, Int $=12.90$), its standout offensive metrics suggest an attacking-focused role, such as a forward or advanced midfielder who both scores and sets up goals.
  • Component 2: Balances moderate offense (Shots $=16.93$, SCA $=40.32$) with robust defensive numbers (Tkl $=55.80$, Blocks $=35.20$, Int $=32.76$, Clr $=60.96$). This dual profile points to two-way or defensive-minded midfielders capable of contributing in attack but primarily excelling at regaining possession and thwarting opponents.
  • Component 3: Characterized by very low values across most attacking and defensive actions (e.g., Goals $=0.05$, Shots $=0.05$, Tkl $=1.01$, Blocks $=0.15$). However, Clearances $=12.51$ is non-negligible. These sparse counts could indicate goalkeepers or rare outliers who make few direct contributions in open play but occasionally clear the ball.
  • Component 4: Dominated by extremely high Clearances (93.07), combined with substantial Tackles (37.63), Blocks (31.98), and Interceptions (29.42). This defensive emphasis overshadows a moderate attacking presence (Shots $=11.76$, SCA $=15.62$). Consequently, this profile aligns with classic center-backs or defense-first players who make frequent clearances and recoveries but still might push forward occasionally.
  • Component 5: Displays considerable Progressive Carries (61.94), Key Passes (31.13), and Tackles (50.16), along with moderate shooting (Shots $=27.44$). Its overall mix suggests players who energetically move the ball forward while contributing in defense. This could fit full-backs or box-to-box midfielders who link offense and defense through carries, passing, and ball recoveries.
  • Component 6: Features very high Assists (35.57) and Crosses into the Penalty Area (56.84), indicating a strong creative and wide-play dimension. It also has a moderate share of Shots (24.10) and Interceptions (17.09). These attributes align with attacking wing-backs or wide midfielders who frequently deliver crosses, set up goals, and offer some defensive coverage.
Overall, the finite mixture model reveals distinct positional tendencies but, unlike the partial membership framework, does not allow for explicit nuanced roles. Each player is assigned to one of these six clusters according to a posterior probability that captures uncertainty rather than genuine mixed roles. Consequently, for players who genuinely balance multiple responsibilities, interpretation can be less direct than in the PM model. A partial membership framework (PM) may offer clearer insight for players whose roles genuinely span multiple clusters.
@@ 3.4 Model Comparison
We compare the partial membership (PM), mixed membership (MM), and finite mixture models on interpretability, computational cost, and alignment with known football roles. All three methods reveal at least one offensive group (strikers), one defensive group (defenders), and one or more midfield/transition groups, a structure consistent with standard football positions (see Fraley and Raftery, 1998; Forgy, 1965, for a broader discussion on cluster interpretability).
A key difference from MM is that the PM model naturally identifies an archetypal (near-pure) player for each cluster, even when we fix the same number of clusters as the other models. For instance, with $K=4$, the MM model yields the highest membership values of only 0.565, 0.686, 0.576, and 0.656 for its four clusters. These do not indicate true “corner” players who fully exemplify one role. Consequently, such clusters can be harder to label consistently, especially in the finite mixture model, where combining roles in a single cluster can obscure a player’s actual multi-role tendencies. The MM model can capture hybrid behaviors at an attribute level but lacks straightforward “archetypal” players who neatly define each cluster.
By contrast, the PM model’s continuous membership vectors, along with near complete memberships for certain archetypal players, make it more intuitive to map each cluster to a well-known role, such as a striker or a goalkeeper. MM does allow attribute-level mixing, but this can complicate the direct labeling of entire clusters. The finite mixture model, with strictly binary (one-cluster) memberships, often ends up blending distinct roles within a single component.
Computationally, we ran each method for $K=2$ through $K=8$ under the same conditions on a MacBook Air (Apple M3, 16GB RAM) with no other processes running. The partial membership (PM) model took about 38,129 seconds ($\sim$10.6 hours), the mixed membership (MM) model 52,797 seconds ($\sim$14.7 hours), and the finite mixture model 26,320 seconds ($\sim$7.3 hours). These times reflect both the dimensionality of each model’s latent variables and the efficiency of our MCMC implementation.
In particular, PM simultaneously samples a continuous membership vector $\bm{\tau}_{i}$ for each player, while MM requires a potentially high-dimensional set of attribute-level latent labels and can thus converge more slowly. The finite mixture model here follows a “one-cluster-per-player” scheme, which substantially reduces the number of latent indicators relative to MM’s attribute-level labels; as a result, it finishes more quickly than the MM model despite having the same overall range of $K$. In practice, exact run times depend on convergence thresholds, initialization, and hyperparameter settings, so these values should be viewed as indicative rather than absolute.
@@ 4 Conclusions And Future Developments
Partial membership models offer analysts significantly greater flexibility compared to traditional model-based clustering or standard distance-based clustering methods. In our study, we tailored the model specifically for count data and applied it to the analysis of Serie A football players during the 2022/2023 season. Our goal was to estimate the various positions players tend to occupy on the field, in addition to their primary positions, by analyzing their playing styles. We based our analysis on a set of 15 count variables recorded during games, carefully selected to cover the key skills pertinent to each player’s role. This application also allowed us to test the model’s reliability by comparing its results with each player’s actual playing position. Furthermore, when compared with mixed membership and mixture models for count data, the partial membership model yielded more realistic and interpretable results. It excelled particularly in capturing archetypal players, whose distinct characteristics define specific clusters. These archetypes not only aid in clarifying each cluster’s definition but also contribute to the model’s overall stability and interpretability. By serving as benchmarks, they assist in better identifying and understanding the roles of players whose classifications might otherwise be ambiguous. While the partial membership model does require considerable computational resources, we believe its potential applications extend beyond sports analysis, offering valuable insights in fields like social sciences, genetics, natural sciences, and textual analysis. It can address some limitations inherent in mixed membership models.
In the realm of sports analytics, our findings could significantly benefit coaches, talent scouts, team managers, and analysts. Utilizing these insights, they can make more informed decisions regarding team strategy, talent acquisition, and statistical research, thereby enhancing both performance and understanding in football.
Looking ahead, two primary areas present intriguing opportunities for enhancing our model. Firstly, addressing the issue of over-dispersion, which is a common challenge in count data, is of considerable interest. Developing methods to accurately assess and incorporate over-dispersion into the model would enable more precise and reliable predictions. Secondly, adding a temporal dimension to the model opens up another avenue for exploration. By incorporating a temporal aspect, the model could provide a dynamic view of clusters and allow for more nuanced analyses. In this direction, one promising line of research would be to extend our approach to multiple football seasons and focus specifically on outfield players, thus offering deeper longitudinal insights while potentially refining the role-based clusters.
@@ Appendix A Application: Washington DC Bike Data
We calculated the number of times bikes are collected from each of the 660 stations and we modelled these counts using a partial membership model, with the intent to explore the interactions between the bikes stations usage, to improve the allocation of the bikes. Partial membership model suits this type of application because the bikes move between the stations along the day, so the usage of the stations could vary and their membership could be partial. While one station may serve predominantly weekday commuters and another may cater mostly to weekend leisure riders, many stations show a blend of both roles (e.g., moderately busy on weekdays, highly used on weekends). By allowing “partial” membership in multiple clusters, our model better reflects these overlapping behaviors than a strictly single-cluster assignment, providing a richer picture of station usage variability. It should be noted that we do not addressed the temporal dependency as the temporal nature of the data would require. Nevertheless, the approaches appear to identify interesting behaviour in the data, and serve to illustrate the usefulness of the method. We run the model over a range of $K=1,\dots,6$. The model with the lowest WAIC is the one with 5 profiles (or components). For a better visualization, in Figure 8 are represented the natural log of the profiles means, while Figure 9 shows the marginal simplices representing stations’ profile membership.